1. A BODY FALLS FREELY FROM REST. FIND

A. THE DISTANCE IT FALLS IN 3 SECOND

B. THE SPEED AFTER FALLING 256 FEET.

C. THE TIME REQUIRED TO ATTAIN A SPEED OF 80FT/SEC.

D. THE TIME IT WILL TAKE TO FALL 900 FEET?



1
Expert's answer
2021-06-01T10:50:13-0400

a)


"h=0.5(32.174)(3)^2=145\ m"


b)


"v^2=2gh\\v^2=2(32.174)(256)\\v=128\frac{ft}{s}"


c)


"80=32.174t\\t=2.49\ s"

d)


"900=0.5(32.174)t^2\\t=7.48\ s"